OpenStudy (anonymous):

2 System of Equation Questions. -Medal Given- :C

OpenStudy (anonymous):

Q1. Which system of equations is represented by the graph? Picture: http://gyazo.com/0df99a021f0ccfc54cbb5726a4906070.png a line is graphed through points negative 6 comma 7 and negative 1 comma 2. It intersects a parabola that opens up at these two points. A. y = x^2 − 6x − 7 x − y = 1 B. y = x^2 − 6x + 7 x + y = −1 C. y = x^2 + 6x − 7 x − y = 1 D. y = x^2 + 6x + 7 x + y = 1

OpenStudy (anonymous):

Q2. Which system of equations is represented by the graph? image: http://gyazo.com/28a33aa151374ed9c230853ed1ef11b0.png rational graph with asymptotes of x equals 4 and y equals negative 1. Linear graph intersects it at negative 2, negative 2 and 5, 5 Image for answer choices - http://gyazo.com/e3019bd9acff2d3c68999feee19144e7.png y = negative x minus 10, over x minus 4 y = x y = x plus 10, over x minus 4 y = −x y = negative x plus ten divided by the quantity x minus four , y = x y = negative x plus ten divided by the quantity x minus four , y = −x

OpenStudy (anonymous):

They are multiple choices, just suck at graphs ._.

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

@Loser66

OpenStudy (anonymous):

Any idea Loser?

OpenStudy (anonymous):

The line is x+y=1 it goes thru (1,0) and (0,1)

OpenStudy (anonymous):

The parabola equation. Since the y-intercept is >0, you can eliminate A and C

OpenStudy (anonymous):

If you replace x=-1 in x^2 + 6 x + 7 you get 2 hence this is the right one

OpenStudy (anonymous):

The answer is D